1. Compute: (show your work)

a. Compute the present values of the following ordinary annuities.

Payment Years Interest Rate Present Value
$678.09 7 13%
$7,968.26 13 6%
$20,322.93 23 4%

b. Compute the future values of the following ordinary annuities.

Payment Years Interest Rate Future Value
$123 13 13%
$4,555 8 8%
$74,484 5 10%
Answer #1
a) Present value
$       2,998.93 =678.09*(1-(1/(1+13%)^7))/13%
$     70,540.48 =7968.26*(1-(1/(1+6%)^13))/6%
$   301,934.55 =20322.93*(1-(1/(1+4%)^23))/4%
b) Future value
$       3,688.12 =123*((1+13%)^13-1)/13%
$     48,449.84 =4555*((1+8%)^8-1)/8%
$   454,732.27 =74484*((1+10%)^5-1)/10%